OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (-2,-3) and has a slope of 1/2

OpenStudy (anonymous):

I need help urgently

OpenStudy (love_to_love_you):

The slope-intercept form is given as y=mx+b where m is the slope and b is the y-intercept when x=0 or at point (0,b). Using the slope-intercept, we can translate it to the standard form given as Ax+By=C where A, B, and C are constants. Step 2. Now we have to find the line with slope m=1/2 going through point (-2,-2). Step 3. Given two points (x1,y1) and (x2,y2), then the slope m is given as y2-y1 m= ----- x2-x1

OpenStudy (love_to_love_you):

Let (x1,y1)=(-2,-3) or x1=-2 and y1=-3. Let other point be (x2,y2)=(x,y) or x2=x and y2=y

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Then what

OpenStudy (love_to_love_you):

Now we're given m=1/2 . Substituting above values and variables in the slope equation m yields the following steps: y2-y1 m= ----- x2-x1 y2-(-3) m= ----- x2-(-2)

OpenStudy (love_to_love_you):

y2-(-3) 1/2= ----- x2-(-2)

OpenStudy (love_to_love_you):

y2+3 1/2= ----- x2+2 Multiply x+2 to both sides to get rid of denominator on right side of equation.

OpenStudy (anonymous):

Ok

OpenStudy (love_to_love_you):

what do you get?

OpenStudy (anonymous):

X-2y=4

OpenStudy (love_to_love_you):

how did you get that?

OpenStudy (anonymous):

Multiplying e sides

OpenStudy (anonymous):

The answers ae 2x+y=4 2x-y=4 X+2y=4 x-2y=4

OpenStudy (anonymous):

Choices

OpenStudy (love_to_love_you):

d is correct. sorry im in 3 questions right now, I forgot to fix the 1/2

OpenStudy (love_to_love_you):

Great job!

OpenStudy (anonymous):

Thanks if u dont have a probkem can u help me with another

OpenStudy (love_to_love_you):

sure :)

OpenStudy (anonymous):

Find the average rate of change for the given function ffrom x=-3 to x=-1 -1 -2 1 2

OpenStudy (anonymous):

There is graph going through -4,0 and 0,4

OpenStudy (love_to_love_you):

ok let me work on it real quick.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

first find the slope.

OpenStudy (anonymous):

First find the slope of the given co-ordinates in the graph, and then make equation out of it.

OpenStudy (love_to_love_you):

@Sthalha1 do you know the slope formula?

OpenStudy (anonymous):

Mx+b

OpenStudy (anonymous):

the formula for finding slope

OpenStudy (anonymous):

because we need slope to form a equation

OpenStudy (anonymous):

slope=m=y2-y1/x2-x1 <--- you have to remember this from now on

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):